OpenStudy (anonymous):

Below are two different functions, f(x) and g(x). What can be determined about their slopes? f(x) Alex earns 1600 dollars in 400 hours. x g(x) 1 −8 5 12 9 32 The function f(x) has a larger slope. The function g(x) has a larger slope. They both have the same slope. The relationship between slopes cannot be determined. Can I please get help with this question? I WILL GIVE MEDAL AND FAN!!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

what is the slope of f(x)?

OpenStudy (anonymous):

I don't know, that's what i'm stuck on

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

slope is the rate of change

jimthompson5910 (jim_thompson5910):

it says "Alex earns 1600 dollars in 400 hours." so the slope is 1600/400 = 4 this means he earns $4 an hour

jimthompson5910 (jim_thompson5910):

what is the slope of g(x)

OpenStudy (anonymous):

1,or 5 or 9?..

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

we have the two points (1,-8) and (5,12)

jimthompson5910 (jim_thompson5910):

What is the slope of the line through (1,-8) and (5,12)

OpenStudy (anonymous):

12-(-8)/5-1 = 20/4 = 5

jimthompson5910 (jim_thompson5910):

the slope of f(x) is 4 the slope of g(x) is 5

OpenStudy (anonymous):

Thank you! I have another question if u don't mind @jim_thompson5910

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (anonymous):

A function is created to represent the balance on a credit card each month. What restrictions would be made to the range? The range would only include integers. The range would only include positive integers. The range would only include negative integers. The range would include all real numbers. I was choosing between B or D but I'm going for D more.. am i wrong? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you have a negative balance on the credit card?

OpenStudy (anonymous):

IDK.. lol i dont know anything about credit cards.. i've never had one, don't ever plan on having one in the future.

jimthompson5910 (jim_thompson5910):

well credit is something you'll need in the future and I'm reading that it is possible to get a negative balance

jimthompson5910 (jim_thompson5910):

so D looks right you can have values that aren't integers for the balance (like $19.02) so that's why B is only partially true

OpenStudy (anonymous):

okayy and I have the answer for this one too and I just wanna make sure

OpenStudy (anonymous):

To which graph does the point (−1, 3) belong? y ≥ 2x + 6 y ≥ 3x − 1<----- MY ANSWER y ≥ −x + 3 y ≥ −2x + 5 @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I see it. I'm graphing each option

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

I am also getting B. Nice job

OpenStudy (anonymous):

Thank you!! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

If i need your help I'll be posting it here for your help if you don't mind :) I just need someone to check the work for me to make sure I have everything right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure I can check it

OpenStudy (anonymous):

LOL it's my bed time ugh! my mom is making me go to bed.. is there any way you can help tomorrow?

OpenStudy (anonymous):

@jim_thompson5910 tomorrow morning maybe?

jimthompson5910 (jim_thompson5910):

I can help tomorrow, yes

jimthompson5910 (jim_thompson5910):

idk about in the morning, but later in the day

OpenStudy (anonymous):

lol okayy thank you so much for your help! very appreciated!

jimthompson5910 (jim_thompson5910):

sure thing